If y varies inversely as x and y = 3, = when x = 6. Find y when x = 36.
Thanks

Answer by fractalier(6550)

The inverse relation is
y = k/x
Now plug in to find k
3 = k/6
k = 18
Now rewrite it with k and new data...
y = 18/x
y = 18/36
y = 1/2